### v1

@@ Abstract
In this paper, we introduce the applications of third-order reduced biquaternion tensors in color video processing. We first develop algorithms for computing the singular value decomposition (SVD) of a third-order reduced biquaternion tensor via a new Ht-product. As theoretical applications, we define the Moore-Penrose inverse of a third-order reduced biquaternion tensor and develop its characterizations. In addition, we discuss the general (or Hermitian) solutions to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ as well as its least-square solution. Finally, we compress the color video by this SVD, and the experimental data shows that our method is faster than the compared scheme.
Key words: Reduced biquaternion tensor; Ht-product; Singular value decomposition; Moore-Penrose inverse.
@@ 1 Introduction
$\mathbb{H}_{c}=\{a_{1}+a_{2}i+a_{3}j+a_{4}k\ |\ i^{2}=-1,\ j^{2}=1,\ ij=ji=k,\ a_{1},\dots,a_{4}\in\mathbb{R}\}.$
@@
Clearly, $k^{2}=-1,\ ki=ik=-j,\ jk=kj=i$. $\mathbb{H}_{c}$ is a four dimensional commutative algebra over $\mathbb{R}$ and contains zero divisors. It is called Segre biquaternion ([14]), commutative quaternion ([9, 12, 13]) or reduced biquaternion ([13, 17]). In this paper, we call it “reduced biquaternion” ( or RB for short).
@@
Reduced biquaternion $\mathbb{H}_{c}$ has numerous applications in many areas. In signal and color image processing, Pei et al. ([12, 13]) show that the operations of the discrete reduced biquaternion Fourier transforms and their corresponding convolutions and correlations are much simpler than the existing implementations of the Hamilton quaternions, and many questions can be performed simultaneously by $\mathbb{H}_{c}$. They use reduced biquaternion matrices to represent the color images, then compress the images by the SVDs of reduced biquaternion matrices (RBSVDs). For comparison, they found that the calculating of RBSVD is faster than SVD of a quaternion matrix (QSVD), and the reconstruction quality is also better than one by QSVD method.
By adopting the special structure of a real representation of a reduced biquaternion matrix, Ding et al. [4] investigated the least-squares special minimal norm solutions to RB matrix equation $AX=B$ and applied the least-squares minimal norm RB solution to the color image restoration.
@@
A tensor can be regarded as a multidimensional array. People use tensors to collect and employ multi-way data. Tensors and tensor decompositions ([3, 6, 7, 8, 10, 15, 19]) have been applied in many areas such as signal processing, computer science, data mining, neuroscience, and more. Studying tensor rank decomposition may go back to the 1960s: the CANDECOMP/PARAFAC (CP) decomposition ([3]) and the Tucker decomposition ([19]). In [6], Kilmer and Martin defined a closed multiplication operation (called t-product) for two real tensors, and considered some important theoretical properties and practical tensor approximation problems. Qin et al. [15] defined a tensor product for third-order quaternion tensors and discussed the Qt-SVD of a third-order quaternion tensor with applications in color video.
Motivated by the works mentioned above, in particular, the applications of reduced biquaternion matrices in color image processing, we introduce the applications of reduced biquaternion tensors in color video compression.
@@
This paper is organized as follows. We first introduce some operations over reduced biquaternion tensors, and then we define the Ht-product and explore some important properties of third-order RB tensors in Section 2. In Section 3, we investigate the Ht-SVD of a third-order RB tensor and develop a method to compute it. Two theoretical applications of Ht-SVD are given in Sections 4 and 5. That is, in Section 4, the Moore-Penrose inverse of a third-order RB tensor is defined and some properties including algebraic expression derived by the Ht-SVD are discussed. In Section 5, we consider the general/Hertmian/least-squares solutions to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. Finally, in Section 6, we apply this Ht-SVD in color video compression. The experimental data shows that our method has an excellent performance comparing with another method.
@@ 2 The Ht-product of third-order tensors over $\mathbb{H}_{c}$
Throughout this paper, Euler script letters $(\mathcal{A,B,\cdots})$ are used to refer to tensors, while the capital letters $(A,B,\cdots)$ represent matrices. The notations $R^{n_{1}\times n_{2}}$ and $R^{n_{1}\times n_{2}\times n_{3}}$ represent the set of all the matrices of dimension $n_{1}\times n_{2}$ and all the third-order tensors of dimension $n_{1}\times n_{2}\times n_{3}$ over algebra $R$ respectively. For a third-order tensor $\mathcal{A}$, we denote the $i$th horizontal, the $j$th lateral and the $k$th frontal slice by $\mathcal{A}(i,:,:),\mathcal{A}(:,j,:)$ and $\mathcal{A}(:,:,k)$ respectively. For simplicity, let $A^{(k)}$ represent the $k$th frontal slice $\mathcal{A}(:,:,k)$.
Recall that an element $q\in\mathbb{H}_{c}$ can be expressed as $q=q_{a}+jq_{b}=q_{(c),1}e_{1}+q_{(c),2}e_{2}$, where $q_{a}=q_{0}+q_{1}i,q_{b}=q_{2}+q_{3}i,$ $q_{(c),1}=q_{a}+q_{b},\ q_{(c),2}=q_{a}-q_{b}\in\mathbb{C}$, and $e_{1}=(1+j)/2,e_{2}=(1-j)/2\in\mathbb{H}_{c}$. Clearly, $e_{1}$ and $e_{2}$ are two special reduced biquaternions satisfying $e_{1}=e_{1}^{2}=\cdots=e_{1}^{n},e_{2}=e_{2}^{2}=\cdots=e_{2}^{n}$ and $e_{1}e_{2}=0$ ([13]).
For any matrix $A\in\mathbb{H}_{c}^{m\times n}$, we can write it in two forms, that is, $A=A_{0}+A_{1}i+A_{2}j+A_{3}k\doteq B_{1}+jB_{2}$ with $A_{0},\dots,A_{3}\in\mathbb{R}^{m\times n},B_{1}=A_{0}+A_{1}i,B_{2}=A_{2}+A_{3}i\in\mathbb{C}^{m\times n}$. Then its conjugate transpose is given by $A^{\ast}=A_{0}^{T}-A_{1}^{T}i+A_{2}^{T}j-A_{3}^{T}k=B_{1}^{\ast}+jB_{2}^{\ast}$. Moreover, $A\in\mathbb{H}^{n\times n}_{c}$ is said to be Hermitian (resp. unitary) if and only if $A=A^{\ast}$ (resp. $AA^{\ast}=A^{\ast}A=I_{n}$).
Furthermore, if we rewrite a $A\in\mathbb{H}_{c}^{m\times n}$ as $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}$, where $A_{(c),1}$ and $A_{(c),2}$ are complex matrices, then the conjugate transpose of $A$ is $A^{\ast}=A_{(c),1}^{\ast}e_{1}+A_{(c),2}^{\ast}e_{2}$. Thus for any $B=B_{(c),1}e_{1}+B_{(c),2}e_{2}\in\mathbb{H}_{c}^{n\times l}$, $AB=A_{(c),1}B_{(c),1}e_{1}+A_{(c),2}B_{(c),2}e_{2}$, and we can use it to prove that $(AB)^{\ast}=B^{\ast}A^{\ast}$.
@@
We will use Vec operator to convert a tensors $\mathcal{A}$ into a block matrix $\textbf{Vec}(\mathcal{A})$:
@@
and the inverse operation Fold to take the block matrix $\textbf{Vec}(\mathcal{A})$ back to a tensor:
@@
Note that by using the normalized Discrete Fourier transformation (DFT) matrix, a complex circulant matrix can be diagonalized ([6], [15]). Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is a third-order complex tensor. Then
@@
where $\otimes$ denotes the Kronecker product, $\widehat{A}^{(i)}$ ($i=1,\cdots,n_{3}$) are the frontal slices of the tensor $\widehat{\mathcal{A}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ which is the result of DFT of $\mathcal{A}$ along the third mode, $F_{n_{3}}$ is unitary and $F_{n_{3}}(i,j)=\frac{1}{\sqrt{n_{3}}}\omega^{(i-1)(j-1)},\ \omega=exp(-\frac{2\pi i}{n_{3}})$.
@@
The DFT of $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ along the third mode is a tensor $\mathcal{\widehat{A}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ satisfying
$\widehat{\mathcal{A}}(i,j,:)=\sqrt{n_{3}}F_{n_{3}}\mathcal{A}(i,j,:).$
We can verify the relation between $\mathcal{A}$ and $\widehat{\mathcal{A}}$ as follows.
@@
(Ht-product) Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. Then $Ht$-product of $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A}\ast_{Ht}\mathcal{B}=\textbf{Fold}((\textbf{circ}(\mathcal{A}_{1,i})+j\textbf{circ}(\mathcal{A}_{j,k}))\textbf{Vec}(\mathcal{B}))\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}},$
where $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}$, $\mathcal{A}_{1,i},\mathcal{A}_{j,k}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}.$
@@
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{C}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}$ with the corresponding DFTs $-^\widehat{\mathcal{A}},\ \widehat{\mathcal{B}}$ and $\widehat{\mathcal{C}}$. Then $\mathcal{A}\ast_{Ht}\mathcal{B}=\mathcal{C}$ if and only if $\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})=\textbf{diag}(\widehat{\mathcal{C}}).$
Proof.  Since $\widehat{\mathcal{A}}$ is the DFT of $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}$ along the third mode, it follows from (3) and $F_{n_{3}}j=jF_{n_{3}}$ that
$\textbf{Vec}(\widehat{\mathcal{A}})$ $=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})(\textbf{Vec}(\mathcal{A}_{1,i})+j\textbf{Vec}(\mathcal{A}_{j,k}))$
$=\textbf{Vec}(\widehat{\mathcal{A}}_{1,i})+j\textbf{Vec}(\widehat{\mathcal{A}}_{j,k}),$
@@
$\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k}).$ (4)
By Definition 2.2 and identities (2) and (3), we have
@@
$=(\textbf{circ}(\mathcal{A}_{1,i})+j\textbf{circ}(\mathcal{A}_{j,k}))\textbf{Vec}(\mathcal{B})$
$=(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k})](F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}(\mathcal{B})$
$=\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k})]\textbf{Vec}(\widehat{\mathcal{B}}).$
@@
$\textbf{Vec}(\widehat{\mathcal{C}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{C})=[\textbf{diag}(\widehat{\mathcal{A}}_{1,i})+j\textbf{diag}(\widehat{\mathcal{A}}_{j,k})]\textbf{Vec}(\widehat{\mathcal{B}}).$
@@
Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$. Using the method in [15], we calculate the conjugate transpose $\mathcal{A}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$ of $\mathcal{A}$ as follows: first conjugately transposing each frontal slice of $\mathcal{A}$ first, and reversing the order of conjugately transposed frontal slices $2$ through $n_{3}$ next.
@@
The conjugate transpose $\mathcal{A}^{\ast}$ of $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,i},\mathcal{A}_{j,k}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is defined as $\mathcal{A}^{\ast}=\mathcal{A}_{1,i}^{\ast}+j\mathcal{A}_{j,k}^{\ast}$ with $\mathcal{A}_{1,i}^{\ast},\mathcal{A}_{j,k}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$.
@@
The inverse of a Segre quaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is unique if it exists, denoted by $\mathcal{B}=\mathcal{A}^{-1}$.
@@
Proof.  It is easy to see that $\textbf{circ}({\mathcal{A}}^{\ast})=(\textbf{circ}({\mathcal{A}}))^{\ast}$ for $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}.$ According to the equality (2), we have
@@
Proof.  Since $\widehat{{\mathcal{A}}^{\ast}}$ is the DFT of ${\mathcal{A}}^{\ast}$ and $F_{n_{3}}j=jF_{n_{3}}$, it follows from (3) that

### v2

@@ Abstract
In this paper, we introduce the applications of third-order reduced biquaternion tensors in color video processing. We first develop an algorithm for computing the singular value decomposition (SVD) of a third-order reduced biquaternion tensor via a new Ht-product. As theoretical applications, we define the Moore-Penrose inverse of a third-order reduced biquaternion tensor and consider its characterizations via its SVD. Using Moore-Penrose inverses, we mainly discuss the general (or Hermitian) solutions to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ as well as its least-squares solutions. Finally, we develop two algorithms and apply them in color video compression and deblurring, both of which perform better than the compared algorithms.
Key words: Reduced biquaternion tensor; Ht-product; Singular value decomposition; Moore-Penrose inverse; Color video processing.
@@ 1 Introduction
$\mathbb{H}_{c}=\{a_{1}+a_{2}\mathbf{i}+a_{3}\mathbf{j}+a_{4}\mathbf{k}\ |\ \mathbf{i}^{2}=-1,\ \mathbf{j}^{2}=1,\ \mathbf{ij}=\mathbf{ji}=\mathbf{k},\ a_{1},\dots,a_{4}\in\mathbb{R}\}.$
@@
Clearly, $\mathbf{k}^{2}=-1,\ \mathbf{ki}=\mathbf{ik}=-\mathbf{j},\ \mathbf{jk}=\mathbf{kj}=\mathbf{i}$. $\mathbb{H}_{c}$ is a four dimensional commutative algebra over $\mathbb{R}$ and contains zero divisors, and it is often called Segre quaternions ([12]), commutative quaternions ([8, 10, 11]) or reduced biquaternions ([11, 14]). In this paper, we call it “reduced biquaternions” ( or RB for short).
@@
Reduced biquaternions $\mathbb{H}_{c}$ has numerous applications in many areas. In signal and color image processing, Pei et al. ([10, 11]) showed that the operations of the discrete reduced biquaternion Fourier transforms and their corresponding convolutions and correlations were much simpler than the existing implementations of the Hamilton quaternions, and many questions could be performed simultaneously by $\mathbb{H}_{c}$. They used reduced biquaternion matrices to represent the color images, and then compressed the images by the SVDs of reduced biquaternion matrices (RBSVDs). For comparison, their proposed algorithm (RBSVDs) was faster and has better reconstruction quality than the compared algorithm of computing the SVD of a quaternion matrix. By adopting the special structure of a real representation of a reduced biquaternion matrix, Ding et al. [3] investigated the least-squares special minimal norm solutions to RB matrix equation $AX=B$ and applied the least-squares minimal norm RB solution to the color image restoration.
@@
A tensor is an extension of the concept of a matrix, which provides a more general and flexible approach to handle and represent complex multidimensional data structures. Tensors and tensor decompositions ([2, 5, 6, 7, 9, 13, 16]) have been applied in many areas such as signal processing, computer science, data mining, neuroscience, etc. Studying tensor rank decomposition may go back to the 1960s: the CANDECOMP/PARAFAC (CP) decomposition ([2]) and the Tucker decomposition ([16]). In [5], Kilmer and Martin defined a closed multiplication operation (called t-product) for two real tensors, and considered some important theoretical properties and practical tensor approximation problems. Qin et al. [13] defined a tensor product for third-order quaternion tensors and discussed the Qt-SVD of a third-order quaternion tensor with applications in color video. We note that a RB tensor also has good capability in representing color video as it does not lose the correlation between the RGB channels. Thus this paper will explore some fundamental theories of RB tensors and their applications, such as utilizing the proposed theories of SVD, Moore-Penrose inverses, the minimal norm least-squares solution to compress or deblur color videos.
@@
This paper is organized as follows. We first introduce some operations over reduced biquaternion tensors, and then we define the Ht-product and explore some important properties of third-order RB tensors in Section 2. In Section 3, we investigate the Ht-SVD of a third-order RB tensor and develop an algorithm to compute it. Two theoretical applications of Ht-SVD are given in Sections 4 and 5. That is, in Section 4, the Moore-Penrose inverse of a third-order RB tensor is defined and some properties including algebraic expression derived by the Ht-SVD are discussed. In Section 5, we mainly consider the general/Hertmian/least-squares solutions/minimal norm least-squares solution to reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. Finally, in Section 6, we apply our results to color video compression and deblurring. The experimental data show that our methods have excellent performances comparing with other methods.
@@ 2 The Ht-product of third-order tensors over $\mathbb{H}_{c}$
Throughout this paper, Euler script letters $(\mathcal{A,B,\cdots})$ are used to refer to tensors, while the capital letters $(A,B,\cdots)$ represent matrices. The notations $\mathbb{R}^{n_{1}\times n_{2}}$ and $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ represent the set of all the matrices of dimension $n_{1}\times n_{2}$ and all the third-order tensors of dimension $n_{1}\times n_{2}\times n_{3}$ over algebra $\mathbb{R}$, respectively. For a third-order tensor $\mathcal{A}$, we denote the $i$th horizontal, the $j$th lateral and the $k$th frontal slice by $\mathcal{A}(i,:,:),\mathcal{A}(:,j,:)$ and $\mathcal{A}(:,:,k)$ respectively. For simplicity, let $A^{(k)}$ represent the $k$th frontal slice $\mathcal{A}(:,:,k)$.
Recall that any element $q=q_{0}+q_{1}\mathbf{i}+q_{2}\mathbf{j}+q_{3}\mathbf{k}\in\mathbb{H}_{c}$ can be expressed as $q=q_{a}+\mathbf{j}q_{b}=q_{(c),1}e_{1}+q_{(c),2}e_{2}$, where $q_{a}=q_{0}+q_{1}\mathbf{i},q_{b}=q_{2}+q_{3}\mathbf{i},$ $q_{(c),1}=q_{a}+q_{b},\ q_{(c),2}=q_{a}-q_{b}$, and $e_{1}=(1+\mathbf{j})/2,e_{2}=(1-\mathbf{j})/2$. Clearly, $e_{1}$ and $e_{2}$ are two special reduced biquaternions satisfying $e_{1}=e_{1}^{2}=\cdots=e_{1}^{n},e_{2}=e_{2}^{2}=\cdots=e_{2}^{n}$ and $e_{1}e_{2}=0$ ([11]). The modulus of $q$ is $|q|=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}}$.
For any matrix $A\in\mathbb{H}_{c}^{m\times n}$, we can write it in two forms, that is, $A=A_{0}+A_{1}\mathbf{i}+A_{2}\mathbf{j}+A_{3}\mathbf{k}=B_{1}+\mathbf{j}B_{2}$ with $A_{0},\dots,A_{3}\in\mathbb{R}^{m\times n},B_{1}=A_{0}+A_{1}\mathbf{i},B_{2}=A_{2}+A_{3}\mathbf{i}\in\mathbb{C}^{m\times n}$. Then its conjugate transpose is given by $A^{\ast}=A_{0}^{T}-A_{1}^{T}\mathbf{i}+A_{2}^{T}\mathbf{j}-A_{3}^{T}\mathbf{k}=B_{1}^{\ast}+\mathbf{j}B_{2}^{\ast}$, where $B_{1}^{\ast},B_{2}^{\ast}$ denote the conjugate transpose of complex matrices $B_{1},B_{2}$. Moreover, $A\in\mathbb{H}^{n\times n}_{c}$ is said to be Hermitian (resp. unitary) if and only if $A=A^{\ast}$ (resp. $AA^{\ast}=A^{\ast}A=I_{n}$). Furthermore, if we rewrite $A\in\mathbb{H}_{c}^{m\times n}$ as $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}$, where $A_{(c),1}=B_{1}+B_{2}$ and $A_{(c),2}=B_{1}-B_{2}$ are complex matrices, then the conjugate transpose of $A$ is $A^{\ast}=A_{(c),1}^{\ast}e_{1}+A_{(c),2}^{\ast}e_{2}$. Thus for any $B=B_{(c),1}e_{1}+B_{(c),2}e_{2}\in\mathbb{H}_{c}^{n\times l}$, $AB=A_{(c),1}B_{(c),1}e_{1}+A_{(c),2}B_{(c),2}e_{2}$, and we can use it to prove that $(AB)^{\ast}=B^{\ast}A^{\ast}$. For $A=(a_{i_{1}i_{2}})\in\mathbb{H}_{c}^{m\times n}$, we define the Frobenius norm of $A$ as $\|A\|_{F}=\sqrt{\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{n}|a_{i_{1}i_{2}}|^{2}}$. Moreover, it can be verified that $\|A\|_{F}^{2}=\mbox{tr}(\Re(A^{\ast}A)),$ where $\Re(\cdot)$ denotes the real part of a reduced biquaternion matrix.
@@
We will use Vec operator to convert a tensor $\mathcal{A}$ into a block matrix $\textbf{Vec}(\mathcal{A})$:
@@
and use the inverse operation Fold to take the block matrix $\textbf{Vec}(\mathcal{A})$ back to the tensor $\mathcal{A}$:
@@
Note that by using the normalized Discrete Fourier transformation (DFT) matrix, a complex circulant matrix can be diagonalized ([5], [13]). Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is a third-order complex tensor, then
@@
where $\otimes$ denotes the Kronecker product, $\widehat{A}^{(i)}$ ($i=1,\cdots,n_{3}$) are the frontal slices of the tensor $\widehat{\mathcal{A}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ which is the result of DFT of $\mathcal{A}$ along the third mode, $F_{n_{3}}$ is an $n_{3}\times n_{3}$ unitary matrix defined by $F_{n_{3}}(s,t)=\frac{1}{\sqrt{n_{3}}}\omega^{(s-1)(t-1)},\ \omega=exp(-\frac{2\pi\mathbf{i}}{n_{3}})$.
@@
For the tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$, the DFT $\widehat{\mathcal{A}}$ of $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ along the third mode is defined as
$\widehat{\mathcal{A}}(s,t,:)=\sqrt{n_{3}}F_{n_{3}}\mathcal{A}(s,t,:).$
As we can see that after the DFT, the resulting vector $\widehat{\mathcal{A}}(s,t,:)$ is still a vector in $\mathbb{H}^{n_{3}}_{c}$. We can verify the relation between $\mathcal{A}$ and $\widehat{\mathcal{A}}$ as follows.
@@
(Ht-product) Let $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$. Then $Ht$-product of $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A}\ast_{Ht}\mathcal{B}=\textbf{Fold}((\textbf{circ}(\mathcal{A}_{1,\mathbf{i}})+\mathbf{j}\textbf{circ}(\mathcal{A}_{\mathbf{j},\mathbf{k}}))\textbf{Vec}(\mathcal{B}))\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$
@@
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}},\ \mathcal{B}\in\mathbb{H}_{c}^{n_{2}\times n_{4}\times n_{3}}$ and $\mathcal{C}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}$ with the corresponding DFTs $\widehat{\mathcal{A}},\ \widehat{\mathcal{B}}$ and $\widehat{\mathcal{C}}$. Then $\mathcal{A}\ast_{Ht}\mathcal{B}=\mathcal{C}$ if and only if $\textbf{diag}(\widehat{\mathcal{A}})\textbf{diag}(\widehat{\mathcal{B}})=\textbf{diag}(\widehat{\mathcal{C}}).$
Proof.  Since $\widehat{\mathcal{A}}$ is the DFT of $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{\mathbf{j},\mathbf{k}}$ along the third mode, it follows from (3) and $F_{n_{3}}\mathbf{j}=\mathbf{j}F_{n_{3}}$ that
$\textbf{Vec}(\widehat{\mathcal{A}})$ $=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})(\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})+\mathbf{j}\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}}))$
$=\textbf{Vec}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{Vec}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}}),$
@@
$\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}}).$ (4)
By Definition 2.2 and the identities (2) and (3), we have
@@
$=(\textbf{circ}(\mathcal{A}_{1,\mathbf{i}})+\mathbf{j}\textbf{circ}(\mathcal{A}_{\mathbf{j},\mathbf{k}}))\textbf{Vec}(\mathcal{B})$
$=(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}})](F_{n_{3}}\otimes I_{n_{2}})\textbf{Vec}(\mathcal{B})$
$=\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{1}})[\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}})]\textbf{Vec}(\widehat{\mathcal{B}}).$
@@
$\textbf{Vec}(\widehat{\mathcal{C}})=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{C})=[\textbf{diag}(\widehat{\mathcal{A}}_{1,\mathbf{i}})+\mathbf{j}\textbf{diag}(\widehat{\mathcal{A}}_{\mathbf{j},\mathbf{k}})]\textbf{Vec}(\widehat{\mathcal{B}}).$
@@
Suppose $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$. Using a similar approach in [13], we calculate the conjugate transpose $\mathcal{A}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$ of $\mathcal{A}$ as follows: conjugately transposing each frontal slice of $\mathcal{A}$ first, and reversing the order of conjugately transposed frontal slices $2$ through $n_{3}$ next.
@@
The conjugate transpose $\mathcal{A}^{\ast}$ of $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{j,k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,\mathbf{i}},\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ is defined as $\mathcal{A}^{\ast}=\mathcal{A}_{1,\mathbf{i}}^{\ast}+j\mathcal{A}_{\mathbf{j},\mathbf{k}}^{\ast}$ with $\mathcal{A}_{1,\mathbf{i}}^{\ast},\mathcal{A}_{\mathbf{j},\mathbf{k}}^{\ast}\in\mathbb{C}^{n_{2}\times n_{1}\times n_{3}}$.
@@
The inverse of a reduced biquaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n\times n\times m}$ is unique if it exists, denoted by $\mathcal{A}^{-1}$.
@@
Proof.  It is easy to see that $\textbf{circ}({\mathcal{A}}^{\ast})=(\textbf{circ}({\mathcal{A}}))^{\ast}$ for $\mathcal{A}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}.$ According to the identity (2), we have
@@
Proof.  Since $\widehat{{\mathcal{A}}^{\ast}}$ is the DFT of ${\mathcal{A}}^{\ast}$ and $F_{n_{3}}\mathbf{j}=\mathbf{j}F_{n_{3}}$, it follows from (3) that